# FitDiff: Robust monocular 3D facial shape and reflectance estimation using Diffusion Models

### v2

@@ Abstract
The remarkable progress in 3D face reconstruction has resulted in high-detail and photorealistic facial representations. Recently, Diffusion Models have revolutionized the capabilities of generative methods by surpassing the performance of GANs. In this work, we present FitDiff, a diffusion-based 3D facial avatar generative model. Leveraging diffusion principles, our model accurately generates relightable facial avatars, utilizing an identity embedding extracted from an “in-the-wild” 2D facial image. The introduced multi-modal diffusion model is the first to concurrently output facial reflectance maps (diffuse and specular albedo and normals) and shapes, showcasing great generalization capabilities. It is solely trained on an annotated subset of a public facial dataset, paired with 3D reconstructions. We revisit the typical 3D facial fitting approach by guiding a reverse diffusion process using perceptual and face recognition losses. Being the first 3D LDM conditioned on face recognition embeddings, FitDiff reconstructs relightable human avatars, that can be used as-is in common rendering engines, starting only from an unconstrained facial image, and achieving state-of-the-art performance. Please see our project page at fitdiff.github.io.
Keywords:
3D Face Reconstruction Face Analysis Diffusion Models
@@ 1 Introduction
Figure 1: Overview of FitDiff, a latent diffusion-based 3D facial generative network. Starting from Gaussian noise, our method generates facial avatars with relightable reflectance and shape, conditioned on an identity embedding. During sampling, a novel guidance algorithm is applied for further control of the resulting identity. $\mathbf{Z}_{T},\mathbf{Z}_{k}$ and $\mathbf{Z}_{k-1}$ are visualized in the actual picture space for illustration purposes.
@@
A fundamental objective of Computer Vision encompasses photorealistic 3D face reconstruction from a single image, which has gained significant attention from the research community over the past few decades. Its numerous applications in computer graphics, virtual reality, and entertainment, include avatar creation, face animation and manipulation [1, 87, 27, 4, 57]. Despite the notable recent progress, accurate replication of personalized facial reconstructions continues to present a challenge. This is primarily due to the inherent ambiguity present in monocular images, as well as the difficulties associated with handling occlusions and capturing substantial variations in lighting conditions and facial expressions. On top of that, captured 3D facial datasets are still relatively small [106], including biases and lacking generalization in parallel with structural limitations of the current facial reconstruction methods.
@@
The 3D Morphable Model (3DMM) was introduced in the seminal work of Blanz and Vetter [6] and uses Principal Component Analysis (PCA) to model the complex facial shapes and textures of 200 facial scans. Since then, there has been tremendous progress in retrieving 3D facial information from a monocular image like large-scale statistical models utilizing hundreds of subjects [8, 68, 50]. Even though most statistical models can accurately reconstruct the facial shape, they do not perform well in capturing the facial texture [8]. More recently, Generative Adversarial Networks (GANs) [31], and particularly the style-based generators [39, 40, 38], have demonstrated effectiveness in capturing intricate frequencies. This success has resulted in numerous subsequent studies combining these methods with 3DMMs [27, 28, 25], where the facial texture is modeled as a UV map. A fundamental challenge inherent in optimization fitting methods is their susceptibility to outliers, requiring to heuristically initialize the GAN’s $\mathbf{z}$ or $\mathbf{w}$ embedding for the back-propagation during inference to avoid instabilities [47]. Additionally, they suffer from problems like unstable training and mode collapse [78, 42].
@@
A solution to the aforementioned problems is the integration of the recent emerging Diffusion Models (DMs) [33]. These architectures have gained a lot of attention lately, because of their current state-of-the-art performance in image synthesis tasks [19]. Taking inspiration from the thermodynamics [82], DMs define a $T$-length Markov Chain by gradually adding normally distributed random noise to the data and learn to predict the input noise for each step $t\in\{1,2,\cdots,T\}$. They synthesize new data by reversing the diffusion process, starting from a normally distributed random noise and progressively denoising it. This methodology has already been applied to synthesize and manipulate facial images [102, 95, 64]. Nevertheless, it is noteworthy to emphasize that most of these methods employ a conditional mechanism reliant on textual descriptions or other auxiliary information [71], thereby directing their attention not exclusively towards the faithful reconstruction of input facial identity and the exploration of the potential of robust identity embeddings. Moreover, Relightify [64] requires partially completed facial UV maps and third-party extracted facial shapes as input, thus being prone these third-party failures.
@@
This study presents FitDiff, the first multi-modal generative model that synthesizes high-fidelity facial avatars given only an input facial image, starting from randomly initialized Gaussian noise. By harnessing the impressive generation capabilities of LDMs [76], we delve into their potential in the field of 3D facial reconstruction. Our approach enables the synthesis of facial avatars by combining facial UV reflectance maps and facial geometry while utilizing an identity embedding layer as a conditioning mechanism. The diffusion process is applied to the concatenation of the latent vectors of the facial texture maps and shape, while a VQGAN AutoEncoder [21] and a 3DMM model (LSFM [8]) are used to decode them. Moreover, we present a novel facial guidance algorithm incorporated into the reverse diffusion process for accurately reconstructing a target facial identity. FitDiff generates high-fidelity facial avatars while achieving a state-of-the-art identity preservation score. The training of our model involves fitting a facial reconstruction network [44] to a manually selected and curated set of images acquired from the CelebA-HQ dataset [37]. The fitting process results in acquiring facial shape and facial reflectance maps. Overall, in this paper:
@@
  • We introduce FitDiff, a multi-modal diffusion-based generative model that jointly produces facial geometry and appearance. The facial appearance consists of facial diffuse albedo, specular albedo, and normal maps, enabling photorealistic rendering.
@@ 2 Related Work
Figure 2: Differences with existing state-of-the-art methods [44, 64]: Prior works rely on multiple separate models, which can fail on challenging inputs (Fig. 10). In contrast, our method uses only a single Latent Diffusion Model for both shape and reflectance textures prediction, achieving simplified architecture, training, and robustness.
@@ 2.1 Face Modeling
A wide variety of works followed the introduction of the emblematic first 3D Morphable Model (3DMM) [6] trained on 200 distinct subjects. Numerous works have since proposed improvements, especially on realistic expressions  [13, 99, 50, 11, 3, 48, 90, 10, 49], large-scale dataset and model releases  [68, 9, 81, 17]. However, those models were unable to capture high-frequency details due to their linear nature. To deal with this, many studies integrated 3DMMs with Deep Neural Networks [88, 92], Mesh Convolutions [60], GANs [27, 28, 26] and VAEs [5, 54, 96, 74]. Additionally, the photorealistic rendering of implicit representation-based methods [58, 65] led to numerous approaches  [24, 35, 100, 23, 62, 57, 4]. However, their applications and editability remain challenging, hence this work focuses on explicit representations. Emphasizing the acquisition of highly detailed facial texture and faithful reflectance maps, Avatarme [43] and AvatarMe++ [45] treat the facial texture as a combination of diffuse albedo, specular albedo and normals UV maps whereas ReflectanceMM[32] models spatially varying BRDF, while trained in low-cost acquired data. In a different vein, the authors of MICA [106] leverage a robust face recognition network [18] with the objective of generating facial shapes. In [97], dense facial landmarks were introduced for better shape reconstruction, whilst a transformer-based facial reconstruction network was introduced in [104]. The aforementioned techniques and more recent methodologies [22, 46, 15, 72], which utilize displacement maps for finer facial shape details, consistently achieve state-of-the-art performance in facial shape competitions such as NoW [79] and REALY [14]. However, many of these approaches either produce facial texture with baked illumination or fail to generate it altogether. Recent works such as  [44, 20, 55] aim to concurrently reconstruct facial shape and texture but still rely on separate dedicated models for each component (Fig. 2). Closer to our work is AlbedoGAN [72], which introduces a single network for acquiring both facial shape and texture. However, the proposed methodology generates maps containing baked illumination and cannot handle wearables, contrary to our approach.
@@ 2.2 Diffusion Models
Inspired by [82, 84, 86, 85], the authors of [19] showed that DMs can perform better than the widely used GAN-based methods in image synthesis tasks. DMs learn to model a data distribution $\mathcal{P}(x)$ by gradually injecting normally distributed noise as a $T$-length Markov Chain and predicting the denoised data in each step. The reverse denoising process can be used as the sampling procedure. The high-quality samples and the more stable training have attracted the attention of the research community leading to a great variety of applications in image generation [73, 34, 76], text-to-image [61, 41], text-to-3D [70], Pointcloud [105, 101, 56] and Mesh [53] generation. One significant limitation of DMs is their sampling speed, which is hindered by their iterative generative nature. Despite the introduction of DDIMs [83], which helped decrease sampling time, a significant amount of computational resources is still necessary for each iteration step. Initial approaches like [93, 80] and then Latent Diffusion Models (LDMs) [76] perform the diffusion process in the latent space, instead of the pixel space, while achieving comparable performance in image synthesis. Extensions of those methods broaden the applications into medicine [69], text-to-audio [52], and video creation [7]. On top of that, some approaches [51, 98] use a pre-trained LDM as priors. Closest to our work, other diffusion-based facial models [102, 95, 64] were presented. The former two employ a coarse-to-fine approach to attain identity information, unlike our single-stage methodology. Furthermore, Rodin [95] uses an implicit representation [16] for facial shape but doesn’t prioritize faithful reconstruction of the input facial identity. On the contrary, FitDiff integrates a fully-controllable 3DMM model, facilitating the generation of high-quality facial avatars that accurately represent the input identity. On the other hand, Relightify [64] uses an unconditional denoising network for filling partially completed facial UV maps extracted from third-party models. Although it may seem similar to our proposed methodology, FitDiff has several advantages over Relightify, which we discuss in detail in Section 6.
Figure 3: Overview of the main phase of our training scheme: During each training iteration, the facial reflectance maps are initially projected onto the latent space and subsequently concatenated to the latent vector $\mathbf{z}_{0}$, to which noise is introduced. Following the estimation of the initial $\bar{\mathbf{z}}_{0}$, perceptual and face recognition losses are imposed upon the estimated initial avatar.
@@ 3 Method
A facial avatar can be formulated as a combination of a mesh $\mathbf{S}$, and texture $\mathbf{T}$ which is defined as a combination of facial reflectance UV maps, namely diffuse albedo ($\mathbf{A}_{D}$), specular albedo ($\mathbf{A}_{S}$) and normals ($\mathbf{N}$). Also, let us denote an “in-the-wild” image containing a face as $\mathbf{I}$, and $\mathbf{V}_{trgt}$ as the corresponding target identity embedding of the appearing face. Given $\mathbf{V}_{trgt}$ as input, FitDiff generates a 3D facial avatar of the same identity as the one in $\mathbf{I}$, alongside with the current scene’s illumination parameters (ambient, diffuse, and specular lighting and lighting direction). An overview of our method is illustrated in Fig. 1, whereas the rest of the section includes a detailed presentation of our method’s architecture (Sec. 3.1), the proposed conditional input (Sec. 3.2), the training scheme (Sec. 3.3), and finally the identity-guidance sampling procedure (Sec. 3.4).
@@ 3.1 Model Architecture
FitDiff is a latent-diffusion based approach [76]. This is motivated by the challenges posed by the large number of parameters and the computational expenses associated with generating multiple meshes and texture images simultaneously. Thus, we represent facial avatars as a latent vector containing latent information about the shape, facial texture, and scene illumination: $\mathbf{z}=\{\mathbf{z}_{tex}|\mathbf{z}_{shp}|\mathbf{z}_{ill}\}\in\mathbb{R}^{4288}$, where $\mathbf{z}_{tex}\in\mathbb{R}^{4096}$ signifies the facial reflectance latent vector, $\mathbf{z}_{shp}\in\mathbb{R}^{183}$ the latent shape vector and $\mathbf{z}_{ill}\in\mathbb{R}^{9}$ the scene illumination parameters. The latent shape vector $\mathbf{z}_{shp}$ can be further separated into the identity parameters $\mathbf{z}_{shp_{i}}\in\mathbb{R}^{158}$ and expression parameters $\mathbf{z}_{shp_{e}}\in\mathbb{R}^{25}$.
@@ 3.2 Conditional input
In the underlying UNet model, we integrate a powerful conditioning mechanism to effectively learn all the necessary identity information. An ideal identity embedding must contain both low and high-frequency information aiming to accurately reconstruct the desired facial avatar. To acquire such an identity embedding, we employ a powerful identity recognition network [18], to which we feed the input facial image. Then, the resulting conditioning vector is a combination of the last feature vector with the intermediate activation layers of the identity recognition network.
@@ 3.3 Model Training
where $\mathbf{A}_{D}$ represents the diffuse albedo, $\mathbf{A}_{S}$ the specular albedo and $\mathbf{N}$ the normals. Our approach includes a branched multi-modal discriminator [44] in combination with a perceptual loss [103] as the training loss. The discriminator is a path-based discriminator [21] comprising two branches, to accommodate their different statistics [44]. The first branch gets as input the concatenation of diffuse and specular albedos $\mathbf{A}_{D}\bigoplus\mathbf{A}_{S}$, whereas the second branch receives the normals $\mathbf{N}$. Furthermore, we adhere to the default training parameters outlined in [76]. For a given triplet $x_{k}$, the encoder $\mathcal{E}$ projects $x_{k}$ into a latent representation $\mathbf{z}_{tex}=\mathcal{E}(x_{k})$, where $\mathbf{z}_{tex}\in\mathbb{R}^{h\times w\times c}$. Then, the latent vector $\mathbf{z}_{tex}$ is fed into the decoder $\mathcal{D}$, which produces a reconstructed output triplet $\bar{x}_{k}$. We downsample the input texture UVs by a factor of $f=H/h=512/64=8$, following the downsampling investigations in  [76] and due to computational limitations. For the shape decoder, we use an LSFM [8] model, pre-trained in $\sim$10k identities.
After training the texture AutoEncoder, we freeze its weight parameters and embark on the training of the identity-conditioned UNet. A comprehensive overview of this training approach is presented in Fig. 3. During this phase, we employ multiple components including the identity embedding $\mathbf{V}_{tgt}$, the shape $\mathbf{z}_{shp}$, the distinct facial texture maps $\mathbf{A}_{D},\mathbf{A}_{S},$ and $\mathbf{N}$, and the scene lighting $\mathbf{z}_{ill}$. At each training step, the facial reflectance maps undergo an initial encoding by the pre-trained encoder $\mathcal{E}$, thereby yielding the latent texture vector $\mathbf{z}_{tex}$. Then, the input vectors are concatenated, i.e., $\mathbf{z}_{0}=\{\mathbf{z}_{tex}|\mathbf{z}_{shp}|\mathbf{z}_{ill}\}$. Let us denote $\mathbf{z}_{t}$ the noisy counterparts of $\mathbf{z}_{0}$, resulting from $t$ steps of noise injection. The UNet network gets $\mathbf{z}_{t}$ as input, and learns to predict the injected noise following:
@@
where $\epsilon$ is the ground-truth injected noise, $\epsilon_{\theta}$ the predicted injected noise from $\mathbf{z}_{t}$, $t$ the diffusion time step, and $\mathbf{V}$ signifies the 2D identity embedding vector as described in Sec. 3.2.
In addition to the primary loss function $\mathcal{L}_{noise}$, our training scheme integrates additional losses intended to enhance robustness, which are the identity losses [27, 44] $\mathcal{L}_{id},\mathcal{L}_{per}$ and the shape loss $\mathcal{L}_{verts}$. Detailed definitions for those are provided in the supplemental. It is important to note that these auxiliary losses are not applicable to the latent variables. Thus, the estimated initial latent vector $\bar{\mathbf{z}}_{0}$ is computed as: $\bar{\mathbf{z}}_{0}=\frac{\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}$. The vector $\bar{\mathbf{z}}_{0}$ is further decoded into the estimated initial avatar. Firstly, $\bar{\mathbf{z}}_{0}$ is split into the estimated initial latent texture vector $\bar{\mathbf{z}}_{0_{tex}}$, latent shape vector $\bar{\mathbf{z}}_{0_{shp}}$ and scene parameters $\bar{\mathbf{z}}_{0_{ill}}$. The first two vectors are fed into the decoder $\mathcal{D}$ and the PCA model $\mathcal{F}_{shp}$ respectively. In this way, the estimated initial facial reflectance maps $\bar{\mathcal{T}}_{0}$ and shape $\bar{\mathcal{S}}_{0}$ are retrieved. Additionally, under the estimated initial scene illumination $\bar{\mathbf{z}}_{0_{ill}}$, we acquire the initial identity rendering $\bar{I}_{0}$ using a differentiable renderer [75], while using a differentiable multi-texture map shader as introduced in  [45, 44] under a single directional light. Overall, the conditional denoising model is trained using the following formula:
@@
Even though both the forward and the reverse diffusion processes can be described via stochastic differential equations in a continuous time [86], they can also be applied in discrete time by choosing a very small step, each time. The selection of the appropriate number of diffusion steps is based on the premise that, in the final step, the input data should be completely converted into random noise. We follow the training parameters proposed by the authors of [76] and we choose $T=1000$ while using a linear noise schedule. Furthermore, FitDiff is trained following the Classifier-Free approach [34]. This means that, during training, we randomly set the input identity embedding equal to zero with a probability of $\mathcal{P}_{uncond}=0.1$.

### v3

@@ Abstract
The remarkable progress in 3D face reconstruction has resulted in high-detail and photorealistic facial representations. Recently, Diffusion Models have revolutionized the capabilities of generative methods by surpassing the performance of GANs. In this work, we present FitDiff, a diffusion-based 3D facial avatar generative model. Leveraging diffusion principles, our model accurately generates relightable facial avatars, utilizing an identity embedding extracted from an “in-the-wild” 2D facial image. The introduced multi-modal diffusion model is the first to concurrently output facial reflectance maps (diffuse and specular albedo and normals) and shapes, showcasing great generalization capabilities. It is solely trained on an annotated subset of a public facial dataset, paired with 3D reconstructions. We revisit the typical 3D facial fitting approach by guiding a reverse diffusion process using perceptual and face recognition losses. Being the first 3D LDM conditioned on face recognition embeddings, FitDiff reconstructs relightable human avatars, that can be used as-is in common rendering engines, starting only from an unconstrained facial image, and achieving state-of-the-art performance.
@@ 1 Introduction
Figure 1: Overview of FitDiff, a diffusion-based 3D facial generative network. Starting from Gaussian noise, our method generates facial avatars with relightable reflectance and shape, conditioned on an identity embedding. During sampling, a novel guidance algorithm ($\mathcal{G}$) is applied for further control of the resulting identity. $\mathbf{Z}_{T},\mathbf{Z}_{k}$ and $\mathbf{Z}_{k-1}$ are visualized in the actual picture space for illustration purposes.
@@
A fundamental objective of Computer Vision encompasses photorealistic 3D face reconstruction from a single image, which has gained significant attention from the research community over the past few decades. Its numerous applications in computer graphics, virtual reality, and entertainment, include avatar creation, face animation and manipulation [1, 83, 26, 4, 54]. Despite the notable recent progress, accurate replication of personalized facial reconstructions continues to present a challenge. This is primarily due to the inherent ambiguity present in monocular images, associated with handling occlusions and capturing substantial variations in lighting conditions and facial expressions. On top of that, captured 3D facial datasets are still relatively small [101], including biases and lacking generalization.
@@
Since the introduction of the first 3D Morphable Model (3DMM) [6], there has been tremendous progress in retrieving 3D facial shape information from a monocular image thanks to large-scale statistical models utilizing hundreds of subjects [7, 65, 49]. More recently, Generative Adversarial Networks (GANs) [30], and particularly the style-based generators [38, 39, 37], have demonstrated effectiveness in capturing intricate facial frequencies resulting into numerous subsequent studies [26, 27, 24]. A fundamental challenge inherent in optimization fitting methods is their susceptibility to outliers, requiring to heuristically initialize the GAN’s $\mathbf{z}$ or $\mathbf{w}$ embedding for the back-propagation during inference to avoid instabilities [46]. Additionally, they suffer from problems like unstable training and mode collapse [74, 41].
@@
A solution to the aforementioned problems is the integration of the recent emerging Diffusion Models (DMs) [32]. Taking inspiration from the thermodynamics [78], DMs define a $T$-length Markov Chain by gradually adding normally distributed noise to the data and learn to predict the input noise for each step $t\in\{1,2,\cdots,T\}$. This methodology has already been applied to synthesize and manipulate facial images [97, 91, 61]. Nevertheless, most of these methods employ a conditional mechanism reliant on textual descriptions or other auxiliary information [67], thereby directing their attention not exclusively towards the faithful reconstruction of the input facial identity and the exploration of the potential of robust identity embeddings. Moreover, Relightify [61] requires partially completed facial UV maps and third-party extracted facial shapes as input, thus being prone to these third-party failures.
@@
This study presents FitDiff, the first multi-modal diffusion model that synthesizes high-fidelity facial avatars given only an input facial image, starting from a randomly initialized Gaussian noise. By harnessing the impressive generation capabilities of LDMs [72], we delve into their potential in 3D facial reconstruction. Our approach enables the synthesis of facial avatars by combining facial UV reflectance maps and facial geometry while utilizing an identity embedding layer as a conditioning mechanism. The diffusion process is applied to the concatenation of the latent vectors of the facial texture maps and shape, while a VQGAN AutoEncoder [20] and a 3DMM model (LSFM [7]) are used to decode them. Moreover, we present a novel facial guidance algorithm incorporated into the reverse diffusion process for accurately reconstructing a target facial identity. FitDiff generates high-fidelity facial avatars while achieving a state-of-the-art identity preservation score. The training of our model involves fitting a facial reconstruction network [43] to a manually selected and curated set of images acquired from the CelebA-HQ dataset [36]. The fitting process results in acquiring facial shape and facial reflectance maps. Overall, in this paper:
@@
  • We introduce FitDiff, a multi-modal diffusion-based generative model that jointly produces facial geometry and appearance. The facial appearance consists of diffuse albedo, specular albedo, and normal maps, enabling photorealistic rendering.
@@ 2 Related Work
Figure 2: Differences with existing state-of-the-art methods [43, 61]: Prior works rely on multiple separate models, which can fail on challenging inputs (Fig. 11). In contrast, our method uses only a single Latent Diffusion Model for both shape and reflectance texture prediction, achieving simplified architecture, training, and robustness.
@@ 2.1 Face Modeling
Various works followed the introduction of the emblematic first 3D Morphable Model (3DMM) [6] trained on 200 distinct subjects. Numerous works have since proposed improvements, especially on realistic expressions  [12, 94, 49, 10, 3, 47, 86, 9, 48], large-scale dataset and model releases  [65, 8, 77, 16]. However, those models could not capture high-frequency details due to their linear nature. To deal with this, many studies integrated 3DMMs with Deep Neural Networks [84, 88], Mesh Convolutions [57], GANs [26, 27, 25] and VAEs [5, 51, 92, 70]. Additionally, the photorealistic rendering of implicit representation-based methods [55, 62] led to numerous approaches  [23, 34, 95, 22, 59, 54, 4]. However, their applications and editability remain challenging, hence this work focuses on explicit representations. Emphasizing the acquisition of highly detailed facial texture and faithful reflectance maps, Avatarme [42] and AvatarMe++ [44] treat the facial texture as a combination of diffuse albedo, specular albedo and normals UV maps whereas ReflectanceMM[31] models spatially varying BRDF, while trained in low-cost acquired data. In a different vein, the authors of MICA [101] leverage a robust face recognition network [17] to generate facial shapes. In [93], dense facial landmarks were introduced for better shape reconstruction, whilst a transformer-based facial reconstruction network was introduced in [99]. The techniques above and more recent methodologies [21, 45, 14, 68], which utilize displacement maps for finer facial shape details, consistently achieve state-of-the-art performance in facial shape competitions such as NoW [75] and REALY [13]. However, many of these approaches produce facial texture with baked illumination or fail to generate it. Recent works such as  [43, 19, 52] aim to concurrently reconstruct facial shape and texture but still rely on separate dedicated models for each component (Fig. 2). Closer to our work is AlbedoGAN [68], which introduces a single network for acquiring both facial shape and texture. However, contrary to our approach, the proposed methodology generates maps containing baked illumination and cannot handle wearables.
@@ 2.2 Diffusion Models
Inspired by [78, 80, 82, 81], the authors of [18] showed that DMs can perform better than the widely used GAN-based methods in image synthesis tasks. The high-quality samples and the more stable training have attracted the attention of the research community leading to a great variety of applications in image generation [89, 76, 72, 69, 33, 72], text-to-image [58, 40], text-to-3D [66], Pointcloud [100, 96, 53] and Mesh [50] generation. Closest to our work, other diffusion-based facial models [97, 91, 61] were presented. Unlike our single-stage methodology, the former two employ a coarse-to-fine approach to attain identity information. Furthermore, Rodin [91] uses an implicit representation [15] for facial shape but doesn’t prioritize faithful reconstruction of the input facial identity. On the contrary, FitDiff integrates a fully controllable 3DMM model, facilitating the generation of high-quality facial avatars that accurately represent the input identity. On the other hand, Relightify [61] uses an unconditional denoising network for filling partially completed facial UV maps extracted from third-party models. Although it may seem similar to our proposed methodology, FitDiff has several advantages over Relightify, which we discuss in detail in Section 6.
@@ 3 Method
A facial avatar can be formulated as a combination of a mesh $\mathbf{S}$ and texture $\mathbf{T}$ which is defined as a combination of facial reflectance UV maps, namely diffuse albedo ($\mathbf{A}_{D}$), specular albedo ($\mathbf{A}_{S}$) and normals ($\mathbf{N}$). Also, let us denote an “in-the-wild” image containing a face as $\mathbf{I}$, and $\mathbf{V}_{trgt}$ as the corresponding target identity embedding of the appearing face. Given $\mathbf{V}_{trgt}$ as input, FitDiff generates a 3D facial avatar of the same identity as the one in $\mathbf{I}$, alongside with the current scene’s illumination parameters (ambient, diffuse, and specular lighting and lighting direction). An overview of our method is illustrated in Fig. 1, whereas the rest of the section includes a detailed presentation of our method’s architecture (Sec. 3.1), the proposed conditional input (Sec. 3.2), the training scheme (Sec. 3.3), and finally the identity-guidance sampling procedure (Sec. 3.4).
@@ 3.1 Model Architecture
FitDiff is a latent-diffusion based approach [72]. This is motivated by the challenges posed by the large number of parameters and the computational expenses associated with simultaneously generating multiple meshes and texture images. Thus, we represent facial avatars as a latent vector containing latent information about the shape, facial texture, and scene illumination: $\mathbf{z}=\{\mathbf{z}_{tex}|\mathbf{z}_{shp}|\mathbf{z}_{ill}\}\in\mathbb{R}^{4288}$, where $\mathbf{z}_{tex}\in\mathbb{R}^{4096}$ signifies the facial reflectance latent vector, $\mathbf{z}_{shp}\in\mathbb{R}^{183}$ the latent shape vector and $\mathbf{z}_{ill}\in\mathbb{R}^{9}$ the scene illumination parameters. The latent shape vector $\mathbf{z}_{shp}$ can be further separated into the identity parameters $\mathbf{z}_{shp_{i}}\in\mathbb{R}^{158}$ and expression parameters $\mathbf{z}_{shp_{e}}\in\mathbb{R}^{25}$.
@@ 3.2 Conditional input
In the underlying UNet model, we integrate a powerful conditioning mechanism to effectively learn all the necessary identity information. An ideal identity embedding must contain low and high-frequency information to accurately reconstruct the desired facial avatar. To acquire such an identity embedding, we employ a powerful identity recognition network [17], to which we feed the input facial image. Then, the resulting conditioning vector combines the last feature vector with the intermediate activation layers of the identity recognition network.
@@ 3.3 Model Training
where $\mathbf{A}_{D}$ represents the diffuse albedo, $\mathbf{A}_{S}$ the specular albedo and $\mathbf{N}$ the normals. Our approach includes a branched multi-modal discriminator [43] in combination with a perceptual loss [98] as the training loss. The discriminator is a path-based discriminator [20] comprising two branches, accommodating their different statistics [43]. The first branch gets as input the concatenation of diffuse and specular albedos $\mathbf{A}_{D}\bigoplus\mathbf{A}_{S}$, whereas the second branch receives the normals $\mathbf{N}$. Furthermore, we adhere to the default training parameters outlined in [72]. For a given triplet $x_{k}$, the encoder $\mathcal{E}$ projects $x_{k}$ into a latent representation $\mathbf{z}_{tex}=\mathcal{E}(x_{k})$, where $\mathbf{z}_{tex}\in\mathbb{R}^{h\times w\times c}$. Then, the latent vector $\mathbf{z}_{tex}$ is fed into the decoder $\mathcal{D}$, which produces a reconstructed output triplet $\bar{x}_{k}$. We downsample the input texture UVs by a factor of $f=H/h=512/64=8$, following the downsampling investigations in  [72] and due to computational limitations. We use the LSFM [7] model for the shape decoder, pre-trained with $\sim$10k identities.
After training the texture AutoEncoder, we freeze its weight parameters and embark on training the identity-conditioned UNet. During this phase, we employ multiple components including the identity embedding $\mathbf{V}_{tgt}$, the shape $\mathbf{z}_{shp}$, the distinct facial texture maps $\mathbf{A}_{D},\mathbf{A}_{S},$ and $\mathbf{N}$, and the scene lighting $\mathbf{z}_{ill}$. At each training step, the facial reflectance maps undergo an initial encoding by the pre-trained encoder $\mathcal{E}$, thereby yielding the latent texture vector $\mathbf{z}_{tex}$. Then, the input vectors are concatenated, i.e., $\mathbf{z}_{0}=\{\mathbf{z}_{tex}|\mathbf{z}_{shp}|\mathbf{z}_{ill}\}$. Let us denote $\mathbf{z}_{t}$ the noisy counterparts of $\mathbf{z}_{0}$, resulting from $t$ steps of noise injection. The UNet network gets $\mathbf{z}_{t}$ as input and learns to predict the injected noise following:
@@
where $\epsilon$ is the ground truth injected noise, $\epsilon_{\theta}$ the predicted injected noise from $\mathbf{z}_{t}$, $t$ the diffusion time step, and $\mathbf{V}$ signifies the 2D identity embedding vector (Sec. 3.2).
In addition to the primary loss function $\mathcal{L}_{noise}$, our training scheme integrates additional losses intended to enhance robustness, which are the identity losses [26, 43] $\mathcal{L}_{id},\mathcal{L}_{per}$ and the shape loss $\mathcal{L}_{verts}$. A detailed definition of those losses is provided in the supplemental material along with more training details. It is important to note that these auxiliary losses do not apply to the latent variables. Thus, in every training step, the initial latent vector $\bar{\mathbf{z}}_{0}$ is estimated by: $\bar{\mathbf{z}}_{0}=\frac{\mathbf{z}_{t}-\sqrt{1-\bar{\alpha}}\epsilon}{\sqrt{\bar{\alpha}_{t}}}$. The vector $\bar{\mathbf{z}}_{0}$ is decoded into the estimated initial avatar. Firstly, $\bar{\mathbf{z}}_{0}$ is split into the estimated initial latent texture vector $\bar{\mathbf{z}}_{0_{tex}}$, latent shape vector $\bar{\mathbf{z}}_{0_{shp}}$ and scene parameters $\bar{\mathbf{z}}_{0_{ill}}$. The first two vectors are fed into the decoder $\mathcal{D}$ and the PCA model $\mathcal{F}_{shp}$ respectively. In this way, the estimated initial facial reflectance maps $\bar{\mathcal{T}}_{0}$ and shape $\bar{\mathcal{S}}_{0}$ are retrieved. Additionally, under the estimated initial scene illumination $\bar{\mathbf{z}}_{0_{ill}}$, we acquire the initial identity rendering $\bar{I}_{0}$ using a differentiable renderer [71] while using a differentiable multi-texture map shader as introduced in  [44, 43] under a single directional light. Overall, the conditional denoising model is trained using the following formula:
@@
Even though both the forward and the reverse diffusion processes can be described via stochastic differential equations in a continuous time [82], they can also be applied in discrete time by choosing a very small step each time. The selection of the appropriate number of diffusion steps is based on the premise that, in the final step, the input data should be completely converted into random noise. We follow the training parameters proposed by the authors of [72], and we choose $T=1000$ while using a linear noise schedule. Furthermore, FitDiff is trained following the Classifier-Free approach [33], aiming to generate novel identities without any prior input. During training and with a probability of $\mathcal{P}_{uncond}=0.1$, we randomly set the input identity embedding equal to zero by feeding the face recognition network [17] with an empty image.